OpenStudy (anonymous):

P= -2x^2 +8x -5 how do you calculate the max. or min. profit of this equation?

OpenStudy (anonymous):

max is at the vertex. first coordinate of the vertex is \(-\frac{b}{2a}\) in your case it is at \(-\frac{8}{2\times (-2)}=2\)

OpenStudy (anonymous):

find the second coordinate by replacing x by 2